# Sharing Matters: Analysing Neurons Across Languages and Tasks in LLMs

### v2

@@ A.3 Prompts
A.4 Supplemental Results of Neuron Behavior Analysis
A.4.1 Neurons Activation Patterns across LLMs
Do the above-mentioned neuron activation patterns change over different LLMs? We further study the neuron activation patterns in another decoder-only multilingual LLM (XGLM) and an encoder-decoder multilingual LLM mT0. The results of XGLM backbones are captured in Figure 12, and the results of mT0-encoder and mT0-decoder are shown in Figure 13 and Figure 14.
The number of active neurons increase in the lower layers, followed by a decrease moving onwards and a rise in the upper layers, despite the absolution values are different from those obtained for BLOOMZ-7b. Moreover, it could be observed that encoder in mT0 involves more all-shared neurons compared to the proportion in decoder.
A.4.2 Neurons Activation Across Examples for Other Tasks
We conduct additional experiments on the XNLI task and Cross-lingual KE (EN (edit) ALL (Test)) task to investigate the type-repeating for neurons. The results shown in Figures 15- 16 are consistent with the results of the Fact Probing task, where only a few neurons retain the same classified types across all examples in the test set.
A.4.3 Neuron Activation Across Languages on XNLI Task
A.4.4 Influence of Model Scale
We investigate neuron activation patterns across the BLOOMZ series with 0.56b, 1b, 3b, 7.1b parameters in a XNLI task. As shown in the results captured in Figure 18, no identifiable pattern difference can be observed to indicate a scale law effect. However, the scale of the model is limited, potentially leading to unreliable results in this experiment. More non-activated neurons in the upper layers of BLOOMZ-7.1b may reflect on a higher level of sparsity for a larger LLM (consistent with Voita et al. (2023); Li et al. (2023)).
A.4.5 Neuron Activation Patterns in Few-shot In-context Learning
According to Wang et al. (2023), in-context learning (ICL) can improve the performance of an LLM under the guidance of few-shot examples in a Cross-lingual KE task. We further explore the impact of few-shot examples on neuron activation patterns. We compare the results of an LLM with 0-shot, 2-shot, 4-shot, 6-shot examples in a Cross-lingual KE (EN (edit) ALL (Test)) task. Four types of neurons in scope have almost identical activation patterns across various few-shot examples (Figure 19). Although in-context examples lead to no observable neuron activation pattern changes, more examples lead to better performances. Could ICL lead to a better neuron activation composition instead of invoking more neurons? We leave this to a future study.
A.4.6 Activation Value across Layers
According to Eq. 1, neurons with larger activation values tend to contribute more to the output. By visualizing neuron activation for a series of thresholds , we can scrutinize the relative importance of various types of activated neurons. The percentage of various types of activated neurons for each activation threshold in the XNLI task are shown in Figure 20. When applying a threshold (>0), there are fewer all-shared, partial-shared, and specific neurons left in the lower layers (layers 0-10) compared to using a lower threshold (i.e., = 0). Under the same threshold scenario, more activated neurons appear in the second half of the model (layers 15-30). It is worth noting these neurons have a higher activation value than neurons activated from a lower threshold (i.e., = 0). Considering both the percentage of activated neurons and their corresponding activation value, it becomes apparent that the neurons in the upper layers contribute more to the output performance.
A.5 Supplemental Results of Neuron Attribution
A.5.1 Contribution Score of Different Tasks
The contribution score of the four types of neurons evaluated on the XNLI task, Cross-lingual KE task, Fact Probing tasks are shown in Figures 24 - 22.
A.5.2 Effective Score of Different Tasks
| all-shared | partial-shared | specific | non-activated | |||||||||
| max | min | mean | max | min | mean | max | min | mean | max | min | mean | |
| en | 1.85 | -0.94 | 0.07 | 0.22 | -0.16 | 0.00 | 0.02 | -0.02 | 0.00 | 0.04 | -0.03 | 0.00 |
| de | 1.03 | -0.60 | 0.02 | 0.13 | -0.13 | 0.00 | 0.07 | -0.03 | 0.00 | 0.02 | -0.01 | 0.00 |
| es | 1.15 | -0.84 | 0.02 | 0.12 | -0.11 | 0.00 | 0.01 | -0.01 | 0.00 | 0.02 | -0.02 | 0.00 |
| fr | 1.06 | -0.78 | 0.01 | 0.15 | -0.11 | 0.00 | 0.03 | -0.04 | 0.00 | 0.02 | -0.02 | 0.00 |
| ru | 0.70 | -0.45 | 0.00 | 0.24 | -0.13 | 0.00 | 0.08 | -0.03 | 0.00 | 0.01 | -0.01 | 0.00 |
| th | 0.50 | -0.90 | 0.00 | 0.17 | -0.10 | 0.00 | 0.03 | -0.05 | 0.00 | 0.01 | -0.01 | 0.00 |
| tr | 0.82 | -0.51 | 0.03 | 0.12 | -0.12 | 0.00 | 0.04 | -0.03 | 0.00 | 0.02 | -0.02 | 0.00 |
| vi | 0.86 | -0.68 | 0.01 | 0.15 | -0.11 | 0.00 | 0.04 | -0.04 | 0.00 | 0.02 | -0.02 | 0.00 |
| zh | 0.52 | -0.42 | 0.00 | 0.17 | -0.20 | 0.00 | 0.08 | -0.07 | 0.00 | 0.02 | -0.01 | 0.00 |
| pt | 1.14 | -0.83 | 0.02 | 0.11 | -0.15 | 0.00 | 0.02 | -0.02 | 0.00 | 0.02 | -0.02 | 0.00 |
The maximum, minimum, and average effective scores of four types of neurons in 10 languages evaluated on the Cross-lingual KE (EN (edit) ALL (Test)) task are shown in Table 3. The maximum effective of four types of neurons across layers evaluated on the Fact Probing task, XNLI task, Cross-lingual KE (ALL (edit) EN (Test)) task are shown in Figures 25 - 27.
A.6 Supplemental Results of Verification on Downstream Tasks
A.6.1 Effects with LLAMA Backbone
| accuracy | pct. | en | de | es | fr | ru | th | tr | vi | zh |
@@ A.6.1 Effects with LLAMA Backbone
| w/o. random | 15% | 52.7 | 44.6 | 47.2 | 46.4 | 44.5 | 38.4 | 40.1 | 48.6 | 45.2 |
| w/o. random | 25% | 46.1 | 42.4 | 41.3 | 43.3 | 40.1 | 34.5 | 39.7 | 38.7 | 40.7 |
| w/o. random | 55% | 28.7 | 30.2 | 28.6 | 30.3 | 25.8 | 19.0 | 27.1 | 28.2 | 25.0 |
In order to further prove the importance of all-shared neurons across LLMs, we conduct the experiments with deactivating neurons on the XNLI task with LLAMA2-7b-chat backbone. The results in Table 4 show that there is more significant effect when all-shared neurons are deactivated. It demonstrates that all-shared neurons play a key role in predicting correct answers across LLMs.

### v3

@@ Appendix C Prompts
Appendix D Supplemental Results on Deactivating Neurons
| settings | pct. | en | de | es | fr | ru | th | tr | vi | zh |
@@ Appendix D Supplemental Results on Deactivating Neurons
| 15% | 52.7 | 44.6 | 47.2 | 46.4 | 44.5 | 38.4 | 40.1 | 48.6 | 45.2 | |
| 25% | 46.1 | 42.4 | 41.3 | 43.3 | 40.1 | 34.5 | 39.7 | 38.7 | 40.7 | |
| 55% | 28.7 | 30.2 | 28.6 | 30.3 | 25.8 | 19.0 | 27.1 | 28.2 | 25.0 |
In order to further prove the importance of all-shared neurons across LLMs, we conduct the experiments with deactivating neurons on the XNLI task with LLAMA2-7b-chat backbone. The results in Table 4 show that there is more significant decline when all-shared neurons are deactivated. It demonstrates that all-shared neurons play a key role in predicting correct answers across LLMs.
Appendix E Generation Impact Score of Different Tasks
The Generation Impact Score of the four types of neurons evaluated on the Cross-lingual KE (EN (edit) ALL (Test)) and XNLI tasks across languages are shown in Figure 7 and Figure 8.
Appendix F Supplemental Results on Neurons Activation Patterns across LLMs
We further study the neuron activation patterns in another multilingual LLM (XGLM). The results of XGLM backbone are captured in Figure 9.
Appendix G Supplemental Results on Neurons Activation Patterns of Foundation LLM BLOOM-7b
We further explore the neuron activation patterns across various tasks in the foundation LLM (BLOOM-7b). The results of BLOOM-7b backbone are captured in Figure 10.
Appendix H Neuron Activation Across Languages on XNLI Task
Appendix I Influence of Model Scale
We investigate neuron activation patterns across the BLOOMZ series with 0.56b, 1b, 3b, 7b parameters in a XNLI task. As shown in the results captured in Figure 12, no identifiable pattern difference can be observed to indicate a scale law effect. However, the scale of the model is limited, potentially leading to unreliable results in this experiment. More non-activated neurons in the upper layers of BLOOMZ-7b may reflect on a higher level of sparsity for a larger LLM.
Appendix J Neuron Activation Patterns in Few-shot In-context Learning
In-context learning (ICL) can improve the performance of an LLM under the guidance of few-shot examples in a Cross-lingual KE task. We further explore the impact of few-shot examples on neuron activation patterns. We compare the results of an LLM with 0-shot, 2-shot, 4-shot, 6-shot examples in a Cross-lingual KE (EN (edit) ALL (Test)) task. Four types of neurons in scope have almost identical activation patterns across various few-shot examples (Figure 13). Although in-context examples lead to no observable neuron activation pattern changes, more examples lead to better performances. Could ICL lead to a better neuron activation composition instead of invoking more neurons? We leave this to a future study.